# Global Compression Commander: Plug-and-Play Inference Acceleration for High-Resolution Large Vision-Language Models

### v4

@@ 10 Additional Experiments
GlobalCom2 3.8 (↓91%) 14.2 (↓40%) 151.0 (↓90%) 13.3 (↓92%) 6.8 (2.7$\times$) 1365.5
Table 8: Efficiency analysis of our GlobalCom2 with LLaVA-NeXT-7B on one NVIDIA A100-SXM4-80GB GPU.
@@
GlobalCom2 7.4 (↓91%) 26.2 (↓29%) 236.0 (↓90%) 25.0 (↓92%) 5.8 (3.1$\times$) 1399.5
Table 9: Efficiency analysis of our GlobalCom2 with LLaVA-NeXT-13B on one NVIDIA A100-SXM4-80GB GPU.
10.4 Detailed Efficiency Analysis.
We conduct a comprehensive analysis of both theoretical and practical efficiency of our GlobalCom2 with LLaVA-NeXT-7B/13B on a single NVIDIA A100-SXM4-80GB GPU, as shown in Table 8 and Table 9. We measure computational efficiency using TFLOPs and peak memory consumption directly, while other metrics are estimated using LLM-Viewer [49]. Given that the sequence length of visual tokens substantially exceeds that of textual and system tokens, we exclude the latter two from our theoretical analysis. For practical efficiency, “Throughout” and the corresponding “Performance” measurements are conducted on the MME benchmark [12], which includes 2374 examples.
As shown in Table 8 and Table 9, GlobalCom2 substantially improves the computational efficiency of LLaVA-NeXT models across both theoretical and practical metrics. Specifically, while preserving model performance, our method achieves significant reductions in GPU memory consumption and notable acceleration in inference speed compared to the original LLaVA-NeXT. Most importantly, these efficiency improvements are achieved through a training-free compression scheme, demonstrating the immediate practical value of GlobalCom2.
10.5 More Visualizations of Token Compression

### v5

@@ E.1 Comparisons on LLaVA-NeXT-13B
[…]
@@
[…]
Table 10: Comprehensive efficiency analysis with LLaVA-NeXT-7B/13B on one NVIDIA A100-SXM4-80GB GPU.
E.3 Sensitivity Analysis of Hyper-parameters
Hyper-parameter $\tau$ serves as the temperature in the softmax function when assessing local crop significance, controlling the “sharpness” of probability distribution. A smaller $\tau$ leads to a sharper distribution, amplifying the differences in global visual importance among local crops. As shown in the first row of Figure 10, our GlobalCom2 demonstrates robust performance across most benchmarks, particularly on VQAT (Singh et al. 2019) and POPE (Li et al. 2023), with varying $\tau$ validating our design principle for HR-LVLMs of allocating retention ratios based on each local crop’s global importance. We empirically set $\tau=10$ to achieve optimal performance across most benchmarks.
Hyper-parameter $\alpha$ determines the criterion for token retention in local crops, where a larger $\alpha$ indicates stronger dependence on global guidance for token retention in local crops. Similar to $\tau$, the second row of Figure 10 shows that our GlobalCom2 exhibits consistent effectiveness across most benchmarks under different $\alpha$ settings, substantiating our dual-perspective token retention strategy of local crops by considering their significance at both global and local levels for HR-LVLMs with dynamic cropping.
E.4 Detailed Efficiency Analysis
We conduct a comprehensive analysis of both theoretical and practical efficiency of our GlobalCom2 with LLaVA-NeXT-7B/13B on a single NVIDIA A100-SXM4-80GB GPU, as shown in Table 10. We measure computational efficiency using TFLOPs and peak memory consumption directly, while other metrics are estimated using LLM-Viewer (Yuan et al. 2024). Given that the sequence length of visual tokens substantially exceeds that of textual and system tokens, we exclude the latter two from our theoretical analysis. For practical efficiency, “Throughout” and the corresponding “Performance” measurements are conducted on the MME (Fu et al. 2023), which includes 2374 examples.
In Table 10, GlobalCom2 substantially improves the computational efficiency of LLaVA-NeXT models across both theoretical and practical metrics. Specifically, while preserving model performance, our method achieves significant reductions in GPU memory consumption and notable acceleration in inference speed compared to the original model. Most importantly, these efficiency improvements are achieved through a training-free compression scheme, demonstrating its practical value.
E.5 More Visualizations of Token Compression